A pair of eyeglass frames are made of an epoxy plastic (coefficient of linear expansion = 1.30 ✕ 10−4°C−1). At room temperature (20.0°C), the frames have circular lens holes 2.37 cm in radius. To what temperature must the frames be heated if lenses 2.38 cm in radius are to be inserted into them?

°C
Answer #1


Chonge in length AL=2.38-2.37= 0.01cm OL=1x104m ZL= LAAT 1x104=2.37x102 x 1:3x10 (Ts -20) 32.46 = T5-20 Ts = 52.46C